Nine pumps can empty a tank in 12 hours how long will it take 8 pumps to empty it?
Thank you,
Marcell

Answer by Alan3354(69443)

Nine pumps can empty a tank in 12 hours how long will it take 8 pumps to empty it?
----------
The job takes 9*12 pump-hours = 108 pump-hr
108 pump-hr/8 pumps = 13.5 hours
-----------
9*12 = 8*t